OpenStudy (anonymous):

I need the answer fast ! medal for reward Find the slope of the line that contains the following points. R(-3, 5), S(3, -2) -7/6 7/6 undefined

OpenStudy (anonymous):

-7/6

OpenStudy (campbell_st):

are you doing a times quiz..?

OpenStudy (anonymous):

ya

OpenStudy (campbell_st):

are you concerned about cheating...?

OpenStudy (anonymous):

\[\frac{ 5--2 }{ -3-3}= -\frac{ 7 }{ 6 }\]

OpenStudy (anonymous):

so it negitive 7

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

The slope is negative because the denominator is negative

OpenStudy (anonymous):

so it would be undefined becausde there is no -6 in my optios

OpenStudy (anonymous):

No, the answer would be -7/6

OpenStudy (anonymous):

oh okay great thanks

OpenStudy (anonymous):

You ARE looking for the slope, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then yes, the answer should be -7/6